One mole of a monoatomic perfect gas is initially at a constant temperature T0 . It expands from a volume V0 to 2V0 under constant pressure. The heat absorbed by the gas is


1
Expert's answer
2020-09-03T14:10:44-0400

The first law of thermodynamic says

"Q=\Delta U+W"

The change in internal energy of one mole of a monoatomic gas

"\Delta U=\frac{3}{2}P(V_2-V_1)\\
=\frac{3}{2}P(2V_0-V_0)=\frac{3}{2}PV_0=\frac{3}{2}RT_0"

The work done by a gas

"W=P(V_2-V_0)=PV_0=RT_0"

Finally

"Q=\frac{3}{2}RT_0+RT_0=\frac{5}{2}RT_0"